From acd@roe.ac.uk Tue Nov 19 09:33:13 2013 Date: Mon, 18 Nov 2013 18:38:34 +0000 From: Clive Davenhall To: rgm@roe.ac.uk, al@roe.ac.uk, mar@roe.ac.uk, etws@roe.ac.uk, msh@roe.ac.uk, njc@roe.ac.uk, nch@roe.ac.uk, rpb@roe.ac.uk, rsc@roe.ac.uk, stv@roe.ac.uk, dmr@roe.ac.uk, acd@roe.ac.uk Cc: dante@astro.puc.cl, jsd@roe.ac.uk, j.p.emerson@qmul.ac.uk, keith@keithnoddle.org, Lorenzo.Rimoldini@unige.ch, mcioni@aip.de, M.J.Jarvis@herts.ac.uk, mike@ast.cam.ac.uk, nigel.metcalfe@durham.ac.uk, norman@astro.gla.ac.uk, pwl@star.herts.ac.uk, rgm@ast.cam.ac.uk, Simon.Dye@nottingham.ac.uk, s.j.warren@ic.ac.uk, tkerr@jach.hawaii.edu, tom.shanks@durham.ac.uk Subject: WFAU VDFS Science Archives progress meeting minutes, 15 November 2013 WFAU VDFS Science Archives progress meeting minutes, 15 November 2013 Present: RGM, NJC, MAR, RPB, ETWS, ACD Apologies: NCH, MSH, KTN, RSC, AL DoNM: 10am Friday 29 November 2013 in the Villas meeting room MAR in the chair (room is booked). Actions discharged this time: ---------------------------- ACTION: RGM will set up a doodle page to arrange the dates for remaining PDRs. Done. However, some members of staff had not yet chosen a date for their PDR, so: **NEW ACTION: All who've not already done so to choose a date for their PDR. ACTION: ACD/RGM to look into maintenance renewals. Done. All pending maintenance renewals had now been renewed. ACTION: MAR/NJC to provide RPB details of the releases to be made public. RPB to do the actual work. Discharged. MAR/NJC had provided RPB with details of the releases, but RPB had not yet done the work. So: **NEW ACTION: RPB to do the work necessary to make the pending releases. There are Viking and VDC releases pending. ACTION NJC: RPB to check the logs, identify the error and forward to NJC, who would investigate and fix it. Discharged This action relates to another VISTA non-survey release. RPB had found and forwarded the logged error to NJC. NJC now needed to fix it. **NEW ACTION: NJC to fix the error. Actions partly discharged but continuing: ---------------------------------------- None. Actions carried forward from 15/11/2013 meeting: ---------------------------------------------- ACTION: ALL to add/maintain entries on http://apache.roe.ac.uk/twiki/bin/view/WFAU/ForthcomingDeadlines Continues. ACTION: MAR to investigate and quantify the issue with unmatched ATLAS sources. And if the issue warrants, RSC to look into the curation code with a view to pairing up the unmatched sources. Continues: MAR had investigated. He had reported his findings to Nigel Metcalfe and Tom Shanks and is still waiting for a response from them before proceeding further. **NEW ACTION: MAR to e-mail Nigel Metcalfe to remind him. ACTION: ACD, RPB to formulate a policy for backups and for checking that data could be recovered from the backups. Continues. ACTION: ACD to discuss with RGM the available budget for hardware purchases. Continues. Progress was being made in understanding the accounts, but was not yet complete. ACTION: NJC sort of the magnitude limits for V1.3 data [sic]. Continues. ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues: this one still has low priority. Specific points and new actions: ------------------------------- Project management: * RGM noted that a statement of interest in the LSST would be submitted shortly. WFCAM, VISTA and VST updates: * RGM noted that he had spoken to Andy Lawrence about interim and final releases of the UKIDSS surveys and Andy, in turn, would speak to the survey leads. Comments and issues arising from CASU minutes: * ETWS noted that the latest CASU minutes suggested that some reprocessing of VISTA data is pending. Networking: * ETWS reported that the networks were performing satisfactorily, but he also noted that better transfer speeds were achieved at night than during the day because the source machine in Cambridge is more heavily loaded during the day. WSA/VSA/OSA/GES Operations: * RPB reported that he had increased the speed of outgesting VVV DR2 data by a factor of two, but that this was inadequate and that a further increase was needed. * NJC suggested that it would be timely to start thinking about restarting processing of DR3 data. * ETWS reported that he had: -- Transferred the latest VISTA data (Aug 2013). -- Ingested the catalogues of the missing OSA data. -- Created metadocs for latest OSA release. ACTION: All; when the OSA release is ready, have a play with accessing it from the new user interface and check that there are no obvious problems. * ACD reported that attempts to ingest the WG10 parameter files into the GES archive were continuing. Various versions of the files had been received, but none would yet ingest cleanly. The WG10 files were the only data still remaining to be ingested. He also noted that Clare Worley and Anna Hourihane would be visiting next week. Hardware and Systems: * MSH had upgraded hardware and software for the following systems: shepseskaf, horus, userkaf and neferefre. These systems now had the latest version of Debian. Subsequently and consequently ETWS had updated their third party software. MAR had performed some rudimentary checks that some of the systems were working adequately. Survey Data Release: * A Viking release was on hold while a number of issues are resolved. * An ATLAS P90 release had gone ahead. However, the release now needs to be re-run to include some missing files that had subsequently come to light. Software: * NJC noted that he was planning more tests for stacks and tiles. * MAR noted that a 2MPZ survey had been prepared as a pseudo-non-survey because password-protected access was required. ACTION NJC to ask Maciej Bilicki to include a note about WFAU hosting the database in the paper describing this survey. * ETWS noted that he had updated code to check ingests for the OSA (incl. missing transfers and unfixed/fixed catalogues). Non-survey Data Release: * MAR reported that a WSA non-survey database had been prepared for Per Friberg. Astrogrid deployment, VO & Data Analysis services: * An internal TAP service had been set up for the latest OSA release, and the associated coverage maps generated and posted. Miscellaneous: * Nothing to report. ACD, draft: 15/10/13, completed: 18/11/13. -- Scanned by iCritical.